# HiTeC: Hierarchical Contrastive Learning on Text-Attributed Hypergraph with Semantic-Aware Augmentation

### v2

@@ Abstract
Contrastive learning (CL) has become a dominant paradigm for self-supervised hypergraph learning, enabling effective training without costly labels. However, node entities in real-world hypergraphs are often associated with rich textual information, which is overlooked in prior works. Directly applying existing CL-based methods to such text-attributed hypergraphs (TAHGs) leads to three key limitations: (1) The common use of graph-agnostic text encoders overlooks the correlations between textual content and hypergraph topology, resulting in suboptimal representations. (2) Their reliance on random data augmentations introduces noise and weakens the contrastive objective. (3) The primary focus on node- and hyperedge-level contrastive signals limits the ability to capture long-range dependencies, which is essential for expressive representation learning. Although HyperBERT pioneers CL on TAHGs, its co-training paradigm suffers from poor scalability. To fill the research gap, we introduce HiTeC, a two-stage hierarchical contrastive learning framework with semantic-aware augmentation for scalable and effective self-supervised learning on TAHGs. In the first stage, we pre-train the text encoder with a structure-aware contrastive objective to overcome the graph-agnostic nature of conventional methods. In the second stage, we introduce two semantic-aware augmentation strategies, including prompt-enhanced text augmentation and semantic-aware hyperedge drop, to facilitate informative view generation. Furthermore, we propose a multi-scale contrastive loss that extends existing objectives with an $s$-walk-based subgraph-level contrast to better capture long-range dependencies. By decoupling text encoder pretraining from hypergraph contrastive learning, this two-stage design enhances scalability without compromising representation quality. Extensive experiments on six real-world datasets validate the effectiveness of our proposed method.
@@ 1 Introduction
Hypergraphs extend traditional graphs by allowing each hyperedge to connect multiple nodes, offering a natural way to model multi-way relationships. Such higher-order structures are prevalent in complex real-world systems, e.g., cortical co-activation in brains (Yu et al. 2011), user-item interactions in e-commerce platforms (Xia, Huang, and Zhang 2022), and multi-author collaborations in academic networks (Kim et al. 2023). In these scenarios, nodes are often associated with rich textual attributes (e.g., product reviews in user–item networks and paper abstracts in co-authorship networks), which provide essential semantic information for downstream applications. While Hypergraph Neural Networks (HNNs) have emerged as a powerful paradigm for hypergraph mining (Chien et al. 2022b; Kim et al. 2024b), effectively capturing high-order structural dependencies, they primarily rely on numerical features and task-specific labels. This reliance limits their effectiveness in real-world applications, where labeled data is scarce and node attributes are predominantly textual. Therefore, underscoring the necessity and significance of self-supervised learning (SSL) on Text-Attributed Hypergraphs (TAHGs).
@@
Recent advances in hypergraph self-supervised learning (HSSL) have primarily focused on contrastive learning paradigms, which aim to maximize the agreement between two augmented views derived from the original hypergraph. HyperGCL (Wei et al. 2022) pioneers this line of work by employing a learnable augmentation strategy. TriCL (Lee and Shin 2023) introduces tri-directional contrastive objectives that model node-, hyperedge-, and membership-level relations in hypergraphs. SE-HSSL (Li et al. 2024) extends this framework by re-designing a hierarchical membership-level objective, achieving state-of-the-art (SOTA) performance. Beyond contrastive methods, generative approaches have also gained initial attention. HypeBoy (Kim et al. 2024a) proposes a hyperedge-filling pretext task to learn expressive hypergraph representations.
@@
Despite recent progress, current HSSL approaches still encounter significant limitations when adopted to hypergraphs with textual attributes: (1) They typically use shallow encoders (e.g., Bag-of-Words (Zhang, Jin, and Zhou 2010) and Skip-Gram (Mikolov et al. 2013)) or pre-trained language models (e.g., BERT (Devlin et al. 2019)) to convert raw text into numerical features, ignoring hypergraph structure. This graph-agnostic design overlooks the correlation between text and higher-order topology, leading to suboptimal representations (Chien et al. 2022a). (2) Current hypergraph contrastive learning (HCL) methods often apply random augmentations to features and structure, such as randomly masking text or hyperlinks. Such perturbations can distort textual meaning and disrupt the alignment between structure and semantic content, thereby introducing noise into the contrastive objective (Fang et al. 2024). (3) Existing HCL methods primarily focus on node- and hyperedge-level signals, while overlooking long-range dependencies and complex topological contexts, which are essential for learning expressive representations (Zhang et al. 2024).
HyperBERT (Bazaga, Lio, and Micklem 2024) takes an early attempt at SSL on TAHGs by jointly encoding textual semantics and hypergraph structure via a Transformer-based architecture. However, its co-training paradigm incurs substantial memory overhead, limiting scalability even on medium-sized hypergraphs (e.g., History and Photo (Yan et al. 2023)), as demonstrated in our experiments. Moreover, it focuses solely on node-level contrastive signals, ignoring multi-scale structural information that has proven beneficial for HSSL (Lee and Shin 2023; Li et al. 2024). Although recent text-attributed graph studies (Chien et al. 2022a; Fang et al. 2024; Zhang et al. 2024) explore joint modeling of text and structure, they overlook high-order relations in hypergraphs, which may limit their applicability to TAHGs.
To address these issues, we propose HiTeC, a hierarchical contrastive learning framework with semantic-aware augmentation for scalable and effective SSL on TAHGs. HiTeC adopts a two-stage design that decouples the pretraining of the text and hypergraph encoders, significantly improving the scalability over joint training paradigms. To overcome the graph-agnostic nature of existing text encoders, we first introduce a structure-aware contrastive objective during text encoder pretraining. Positive and negative text pairs are constructed based on structural proximity, encouraging the encoder to capture correlations between textual content and higher-order topology. Furthermore, to mitigate the noise introduced by random augmentations, we propose two semantic-aware augmentation strategies: a prompt-enhanced text augmentation module, which injects structural context into raw text to generate diverse yet semantically consistent views, and a semantic-aware hyperedge drop mechanism, which discards hyperedges according to their internal semantic cohesion to preserve meaningful relational patterns. Finally, to capture long-range dependencies and multi-scale structural knowledge, we formulate hierarchical contrastive objectives at the node, hyperedge, and subgraph levels. Specifically, our subgraph contrastive objective leverages an $s$-walk-based sampling strategy to extract structurally informative subgraphs, facilitating representation learning over broader contexts. Our main contributions are summarized as follows:
@@
  • We propose HiTeC, a scalable and effective two-stage contrastive learning framework for TAHGs, addressing key limitations of existing HSSL methods in modeling complex structural dependencies and rich textual content.
@@
  • We introduce a structure-aware contrastive objective to pre-train the text encoder, enabling integration of textual content with hypergraph structure, and overcoming the graph-agnostic nature of existing text encoders.
@@
  • We enhance the hypergraph encoder with semantic-aware augmentations and hierarchical contrastive objectives. Specifically, a prompt-enhanced text augmentation and semantic-aware hyperedge drop strategies are designed to reduce perturbation noise. And a novel $s$-walk-based subgraph-level signal is employed to capture long-range structural dependencies.
@@
  • We conduct a comprehensive evaluation of HiTeC on six real-world datasets. The experimental results demonstrate its superiority over strong baselines across various benchmark downstream tasks.
@@ 2 Preliminary
Notations. We define a text-attributed hypergraph (TAHG) as $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{|\mathcal{V}|}$ denotes the set of nodes, and $\mathcal{E}=\{e_{j}\}_{j=1}^{|\mathcal{E}|}$ denotes the set of hyperedges, with each hyperedge $e\in\mathcal{E}$ is a non-empty subset of nodes, and $|e|$ is the size of $e$. Each node $v_{i}$ is associated with a textual sequence $t_{i}\in\mathcal{T}$, and $\mathcal{T}$ represents the set of all text attributes. Each hyperedge $e_{j}$ is associated with a positive scalar weight $w_{j}$, and all weights formulate a diagonal matrix $\mathbf{W}\in\mathbb{R}^{|\mathcal{E}|\times|\mathcal{E}|}$. The hypergraph structure can be modeled as an incidence matrix $\mathbf{H}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}|}$, where $h_{ij}=1$ if node $v_{i}$ belongs to hyperedge $e_{j}$, and $h_{ij}=0$ otherwise. We define the node degree matrix $\mathbf{D}_{v}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, where each entry is $d(v_{i})=\sum_{e_{j}\in\mathcal{E}}w_{j}\cdot h_{ij}$. Similarly, the hyperedge degree matrix is defined as $\mathbf{D}_{e}\in\mathbb{R}^{|\mathcal{E}|\times|\mathcal{E}|}$, where each element $\delta(e_{j})=\sum_{v_{i}\in e_{j}}h_{ij}$ represents the number of nodes connected by the hyperedge $e_{j}$.
Hypergraph neural networks. Hypergraph neural networks (HNNs) have emerged as effective models for representation learning on hypergraphs (Feng et al. 2019; Bai, Zhang, and Torr 2021; Chien et al. 2022b). Most HNNs adopt a two-stage message passing mechanism that alternates between node-to-hyperedge and hyperedge-to-node aggregation. At each layer $l$, the embedding of a hyperedge is updated based on its incident nodes, while a node embedding is updated via its associated hyperedges:
@@
where $z^{(l)}_{e_{j}}$ and $z^{(l)}_{v_{i}}$ denote the embeddings of hyperedge $e_{j}$ and node $v_{i}$ at layer $l$, respectively. The functions $f^{(l)}_{\mathcal{V}\rightarrow\mathcal{E}}$ and $f^{(l)}_{\mathcal{E}\rightarrow\mathcal{V}}$ are permutation-invariant aggregation functions that propagate messages between nodes and hyperedges. This iterative process allows the model to capture complex high-order relations in the hypergraph structure.
Problem statement. Given a TAHG $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, the objective is to learn an optimal function $\Psi:\mathcal{H}\rightarrow(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}})$ in an unsupervised manner that maps the hypergraph to low-dimensional embeddings of nodes and hyperedges. These representations may then be applied to downstream tasks such as node classification and hyperedge prediction.
@@ 3 Methodology
3.1 Overview
As illustrated in Figure 1, we present HiTeC, a contrastive learning framework tailored for TAHGs. In the first stage, we pre-train the text encoder $f_{\theta}$ with a structure-aware contrastive objective, allowing it to capture topological knowledge and achieve more expressive text representations than graph-agnostic encoders. In the second stage, we pre-train a hypegraph encoder $g_{\phi}$ with $f_{\theta}$ frozen. To fully exploit the rich semantic context of nodes, we introduce two semantic-aware augmentation strategies on both text- and structure-level, enhancing view diversity while preserving semantic coherence. Furthermore, we design a hierarchical contrastive objective across node, hyperedge, and subgraph levels. Specifically, the subgraph-level objective, enabled by $s$-walk sampling, captures long-range dependencies. By decoupling the training of two encoders, HiTeC achieves better scalability and performance in integrating textual content and high-order structures.
@@ 3.2 Structure-aware Text Encoder Pretraining
Text encoder. Our text encoder $f_{\theta}$ is built upon a pre-trained language model (PLM) (e.g., BERT (Devlin et al. 2019)) with a lightweight projection layer. To enable efficient training, we apply the parameter-efficient approach, LoRA (Hu et al. 2022). Given a node $v_{i}\in\mathcal{V}$ with textual attribute $t_{i}$, the textual representation is formulated as:
$x_{i}=f_{\theta}(t_{i}),$ (2)
where $x_{i}\in\mathbb{R}^{d}$ denotes the embedding corresponding to the [CLS] token in the output sequence.
Structure-aware contrastive objective. In real-world hypergraphs, nodes are often semantically closer to their local neighbors than to distant ones. For example, in co-citation hypergraphs, papers within the same hyperedge tend to be more topically related. To encode these structural priors, we pre-train $f_{\theta}$ by contrasting local and distant semantic contexts. Specifically, for each node $v_{i}\in\mathcal{V}$, we treat its 1-hop neighbors $N^{(1)}_{v_{i}}$ as positives, and nodes beyond 1-hop distance $\mathcal{V}\setminus N^{(1)}_{v_{i}}$ as negatives. As shown in Figure 1, 1-hop neighbors are defined as nodes belong to the same hyperedges that contain $v_{i}$. Contrastive pairs are constructed by mean pooling embeddings from each set:
$\begin{split}x_{i}^{+}&=\frac{1}{|N^{(1)}_{v_{i}}|}\sum_{v_{j}\in N^{(1)}_{v_{i}}}f_{\theta}(t_{j}),\\ x_{i}^{-}&=\frac{1}{|\mathcal{V}\setminus N^{(1)}_{v_{i}}|}\sum_{v_{j}\in\mathcal{V}\setminus N^{(1)}_{v_{i}}}f_{\theta}(t_{j}),\end{split}$ (3)
where $x_{i}^{+}\in\mathbb{R}^{d}$ and $x_{i}^{-}\in\mathbb{R}^{d}$ denote the aggregated representations of the positive and negative samples, respectively.
A triplet margin loss (Schroff, Kalenichenko, and Philbin 2015) is adopted to guide $f_{\theta}$ in distinguishing positive sample $x_{i}^{+}$ from negative sample $x_{i}^{-}$. For each anchor $x_{i}$, the loss of the text encoder is defined as:
$\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-})=\max\{\cos(x_{i},x_{i}^{+})-\cos(x_{i},x_{i}^{-})+m\},$ (4)
where $\cos(\cdot)$ denotes cosine similarity, and $m$ is the margin hyperparameter. The overall structure-aware objective is averaged across all nodes:
$\mathcal{L}_{t}=\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-}),$ (5)
This process allows $f_{\theta}$ to incorporate structural signals into textual representations.
Figure 2: An example of the prompt template on Cora.

### v3

@@ Abstract
Contrastive learning (CL) has become a dominant paradigm for self-supervised hypergraph learning, enabling effective training without costly labels. However, node entities in real-world hypergraphs are often associated with rich textual information, which has been largely ignored in prior works. Directly applying existing CL-based methods to such text-attributed hypergraphs (TAHGs) leads to three key limitations: (1) The common use of graph-agnostic text encoders fails to capture the correlations between textual semantics and hypergraph topology, resulting in less expressive representations. (2) Their reliance on random data augmentations introduces noise and weakens the contrastive signals. (3) The primary focus on node- and hyperedge-level contrastive signals limits the ability to capture long-range dependencies, which is essential for effective representation learning. To address these challenges, we introduce HiTeC, a two-stage hierarchical contrastive learning framework for effective self-supervised learning on TAHGs. In the first stage, we pre-train the text encoder with a structure-aware contrastive objective to overcome the graph-agnostic nature of conventional methods. In the second stage, we begin by introducing semantic-aware augmentations, including structure-contextualized text augmentation and semantic-aware hyperedge dropping, to facilitate informative view generation. Subsequently, we propose a multi-scale contrastive loss with an $s$-walk-based subgraph-level objective to capture long-range dependencies. Extensive experiments on six real-world datasets validate the effectiveness of our proposed method.
HiTeC: Hierarchical Contrastive Learning on Text-Attributed Hypergraph with Semantic-Aware Augmentation
Mengting Pan1, Fan Li1, Chen Chen2, Xiaoyang Wang1, Wenjie Zhang1 1The University of New South Wales 2University of Wollongong {mengting.pan,fan.li8,xiaoyang.wang1,wenjie.zhang}@unsw.edu.au chenc@uow.edu.au
@@ 1 Introduction
Hypergraphs extend traditional graphs by allowing each hyperedge to connect multiple nodes, offering a natural way to model multi-way relationships. Such higher-order structures are prevalent in complex real-world systems, e.g., cortical co-activation in brains Yu et al. (2011), user-item interactions in e-commerce platforms Xia et al. (2022), and multi-author collaborations in academic networks Kim et al. (2023). In these scenarios, nodes are often associated with rich textual attributes (e.g., product reviews in user–item networks and paper abstracts in co-authorship networks), which provide essential semantic information for downstream applications. While Hypergraph Neural Networks (HNNs) have emerged as a powerful paradigm for hypergraph learning Chien et al. (2022b); Kim et al. (2024b), effectively capturing high-order structural dependencies, they primarily rely on numerical features and task-specific labels. This reliance limits their effectiveness in real-world applications, where labeled data is scarce, and node attributes are predominantly textual. This motivates the development of self-supervised learning (SSL) methods tailored to text-attributed hypergraphs (TAHGs) Bazaga et al. (2024); Nakajima and Uno (2025).
@@
Recent advances in hypergraph self-supervised learning (HSSL) primarily focus on the contrastive paradigm, which maximizes agreement between two augmented views derived from the original hypergraph. HyperGCL Wei et al. (2022) pioneers this direction with a learnable augmentation strategy. TriCL Lee and Shin (2023) introduces tri-directional contrastive objectives over node-, hyperedge-, and membership-level relations in hypergraphs. SE-HSSL Li et al. (2024) further redesigns a hierarchical membership-level objective, achieving state-of-the-art (SOTA) performance. Beyond contrastive methods, generative approaches have also gained initial attention. HypeBoy Kim et al. (2024a) proposes a hyperedge-filling pretext task to learn expressive hypergraph representations.
@@
Despite recent progress, current HSSL approaches still face significant limitations on hypergraphs with textual attributes: (1) They typically use shallow encoders (e.g., Bag-of-Words Zhang et al. (2010) and Skip-Gram Mikolov et al. (2013)) or pre-trained language models (e.g., BERT Devlin et al. (2019)) to convert raw text into numerical features while ignoring hypergraph structure. This graph-agnostic design overlooks the correlation between text and higher-order topology, leading to suboptimal representations Chien et al. (2022a). (2) Current hypergraph contrastive learning (HCL) methods often apply random augmentations to features and structure, e.g., random masking of text or hyperlinks. These perturbations may distort textual meaning and disrupt the alignment between structure and semantic content, introducing noise into the contrastive objective Fang et al. (2024). (3) Existing HCL methods mainly emphasize node- and hyperedge-level signals, but struggle to capture long-range dependencies and richer topological contexts that are crucial for expressive representation learning Zhang et al. (2024).
HyperBERT Bazaga et al. (2024) takes an early attempt at SSL on TAHGs by jointly training text and hypergraph encoders through contrastive learning. However, its contrastive objective mainly relies on 1-hop hyperedge relations, limiting its ability to capture multi-scale dependencies that have proven beneficial for HSSL Lee and Shin (2023); Li et al. (2024). Meanwhile, recent text-attributed graph studies Chien et al. (2022a); Fang et al. (2024); Zhang et al. (2024) explore joint text–structure modeling, but their extension to TAHGs remains challenging due to insufficient modeling of higher-order relations inherent to hypergraphs.
To address these issues, we propose HiTeC, a hierarchical contrastive learning framework for effective SSL on TAHGs. HiTeC adopts a two-stage framework that first injects higher-order topology into text representations and then performs semantic-aware hypergraph contrastive learning. In the first stage, we pre-train a text encoder with a structure-aware contrastive objective to align textual semantics with higher-order topology, producing topology-enhanced text representations. In the second stage, we pre-train a hypergraph encoder with semantic-aware augmentations and hierarchical contrastive objectives. Specifically, we introduce structure-contextualized text augmentation and semantic-aware hyperedge dropping to alleviate perturbation noise. Based on the augmented views, we further formulate node-, hyperedge-, and subgraph-level contrastive objectives, where an $s$-walk-based subgraph signal captures long-range dependencies and richer structural contexts. Our main contributions are as follows:
@@
  • We propose HiTeC, an effective two-stage contrastive learning framework for TAHGs that addresses key limitations of existing HSSL methods in modeling rich textual attributes and complex hypergraph structure.
@@
  • We introduce a structure-aware contrastive objective for text encoder pretraining to integrate textual content with hypergraph structure.
@@
  • We enhance the hypergraph encoder with two semantic-aware augmentation strategies and hierarchical contrastive objectives, where structure-contextualized text augmentation and semantic-aware hyperedge dropping are used to reduce perturbation noise, and an $s$-walk-based subgraph-level contrastive signal captures long-range structural dependencies.
@@
  • Extensive experiments on six real-world datasets demonstrate the effectiveness and superiority of HiTeC over strong baselines across multiple downstream tasks.
@@ 2 Preliminary
Notations. We define a text-attributed hypergraph (TAHG) as $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, where $\mathcal{V}=\{v_{i}\}_{i=1}^{|\mathcal{V}|}$ denotes the set of nodes, and $\mathcal{E}=\{e_{j}\}_{j=1}^{|\mathcal{E}|}$ denotes the set of hyperedges, where each hyperedge $e\in\mathcal{E}$ is a non-empty subset of nodes, and $|e|$ is the size of $e$. Each node $v_{i}$ is associated with a textual sequence $t_{i}\in\mathcal{T}$, where $\mathcal{T}$ denotes the set of all text attributes. Each hyperedge $e_{j}$ is associated with a positive scalar weight $w_{j}$, and all weights formulate a diagonal matrix $\mathbf{W}\in\mathbb{R}^{|\mathcal{E}|\times|\mathcal{E}|}$. The hypergraph structure can be modeled as an incidence matrix $\mathbf{H}\in\{0,1\}^{|\mathcal{V}|\times|\mathcal{E}|}$, where $h_{ij}=1$ if node $v_{i}$ belongs to hyperedge $e_{j}$, and $h_{ij}=0$ otherwise. We define the node degree matrix $\mathbf{D}_{v}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|}$, where each entry is $d(v_{i})=\sum_{e_{j}\in\mathcal{E}}w_{j}\cdot h_{ij}$. Similarly, the hyperedge degree matrix is defined as $\mathbf{D}_{e}\in\mathbb{R}^{|\mathcal{E}|\times|\mathcal{E}|}$, where each element $\delta(e_{j})=\sum_{v_{i}\in e_{j}}h_{ij}$ represents the number of nodes connected by the hyperedge $e_{j}$.
Hypergraph neural networks. HNNs have emerged as effective tools for representation learning on hypergraphs Feng et al. (2019a); Bai et al. (2021); Chien et al. (2022b). Most HNNs adopt a two-stage message passing mechanism that alternates between node-to-hyperedge and hyperedge-to-node aggregation. At each layer $l$, the embedding of a hyperedge is updated based on its incident nodes, while a node embedding is updated via its associated hyperedges:
@@
where $z^{(l)}_{e_{j}}$ and $z^{(l)}_{v_{i}}$ denote the embeddings of hyperedge $e_{j}$ and node $v_{i}$ at layer $l$, respectively. The functions $f^{(l)}_{\mathcal{V}\rightarrow\mathcal{E}}$ and $f^{(l)}_{\mathcal{E}\rightarrow\mathcal{V}}$ are permutation-invariant aggregation functions that propagate messages between nodes and hyperedges. This iterative process allows the model to capture complex higher-order relations in the hypergraph structure.
Problem statement. Given a TAHG $\mathcal{H}=\{\mathcal{V},\mathcal{E},\mathcal{T}\}$, the objective is to learn an optimal function $\Psi:\mathcal{H}\rightarrow(\mathbf{Z}_{\mathcal{V}},\mathbf{Z}_{\mathcal{E}})$ in an unsupervised manner, mapping the hypergraph to low-dimensional node and hyperedge embeddings. These representations may then be applied to downstream tasks such as node classification and hyperedge prediction.
@@ 3 Methodology
In this section, we present HiTeC, a novel and effective SSL framework tailored for TAHGs. As illustrated in Figure 1, HiTeC adopts a two-stage paradigm. In the first stage, a text encoder $f_{\theta}$ is pre-trained with a structure-aware contrastive objective to capture topology-enhanced text representations. In the second stage, with $f_{\theta}$ frozen, we pre-train a hypergraph encoder $g_{\phi}$ to fully exploit the rich semantic context of nodes through semantic-aware augmentations and hierarchical contrastive objectives at the node, hyperedge, and subgraph levels.
@@ 3.1 Structure-aware Text Encoder Pretraining
Text encoder. Our text encoder $f_{\theta}$ is built upon a pre-trained language model, followed by a two-layer MLP projection head that maps the [CLS] representation to a $d$-dimensional embedding. Given a node $v_{i}$ with textual attribute $t_{i}$, the resulting text embedding $x_{i}\in\mathbb{R}^{d}$ is computed as:
$x_{i}=f_{\theta}(t_{i}).$ (2)
Structure-aware contrastive objective. In real-world hypergraphs, nodes are often semantically closer to their local neighbors than to distant ones. To encode this structural prior, we treat the 1-hop neighbors $N^{(1)}_{v_{i}}$ of $v_{i}$, i.e., nodes sharing at least one hyperedge with $v_{i}$, as positives, while the remaining nodes $\mathcal{V}\setminus N^{(1)}_{v_{i}}$ are treated as negatives. Positive and negative representations are constructed by mean pooling the corresponding text embeddings:
$\begin{split}x_{i}^{+}&=\frac{1}{|N^{(1)}_{v_{i}}|}\sum_{v_{j}\in N^{(1)}_{v_{i}}}x_{j},\\ x_{i}^{-}&=\frac{1}{|\mathcal{V}\setminus N^{(1)}_{v_{i}}|}\sum_{v_{j}\in\mathcal{V}\setminus N^{(1)}_{v_{i}}}x_{j}.\end{split}$ (3)
A triplet margin loss Schroff et al. (2015) is adopted to guide $f_{\theta}$ in distinguishing positive sample $x_{i}^{+}$ from negative sample $x_{i}^{-}$. For each anchor $x_{i}$, the loss is defined as:
$\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-})=\max\big(\cos(x_{i},x_{i}^{-})-\cos(x_{i},x_{i}^{+})+m,0\big),$ (4)
where $\cos(\cdot)$ denotes cosine similarity, and $m$ is the margin hyperparameter. The final objective is:
$\mathcal{L}_{t}=\frac{1}{|\mathcal{V}|}\sum_{v_{i}\in\mathcal{V}}\ell_{t}(x_{i},x_{i}^{+},x_{i}^{-}).$ (5)